two planes leave chicago at 8am, one traveling west at a speed 75 mph faster than the other, which is headed east. at 11am, the planes are 2,175 miles apart. how fast are the planes travelling?
Answer by Paul(988)

from 8am to 11am the time difference is 3 hours.
The speed of one plane is 75+x
The speed of another plane is x
Equation:
3(x+75+x)=2175
6x=2175-225
6x=1950
x=325
325+75=400
Hence, the speed of the two planes are 325mph and 400mph.
Paul.